[figure: 9.7 cm]

MAGNETIC PROPERTIES OF LAYER-TYPE COMPOUNDS

TlGdS_{2} AND TlGdSe_{2}

M. Duczmal^{a}, L. Pawlak^{a} and S. Pokrzywnicki^{b} aWrocław University of Technology, Department of Chemistry, Wyb. Wyspianskiego 27, 50-370 Wrocław, Poland

bTechnical University of Opole, Institute of Mathematics, Physics and Chemistry, ul. Luboszycka 7, 45-036 Opole, Poland

Ternary thallium lanthanide dichalcogenides TlLnX_{2} (X = S, Se or Te; Ln = lanthanide, except Ce and Pr) crystallize in the rhombohedral structure of α-NaFeO_{2} type (R¯3m). Their crystal lattice consists of the layers of Ln^{3+} ions separated by three layers of the non-magnetic ions (-Ln-X-Tl-X-Ln-). The magnetization was measured in the field range 0-14 T. The low temperature magnetic susceptibility was also measured (down to 2.2 K for TlGdS_{2}) in order to find a possible magnetic transition. The molecular field constants λ were estimated by fitting the Brillouin function to the experimental magnetization plots. The difference between the λ values for the thallium gadolinium sulphide and the selenide corresponds to the different character of Gd-S and Gd-Se bonds and gives rise to the different J_{1} and J_{2} exchange integrals.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Stefan Pokrzywnicki

Address for correspondence : Politechnika Opolska Instytut Matematyki, Fizyki i Chemii ul. Luboszycka 7 45-036 Opole

Email address : pok@po.opole.pl